# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v1

@@ Appendix A Appendix Summary
Finetuning StrategiesAppendix B): We elaborate on our design choice to freeze the Q-Former in InstructBLIP-based models. This decision is motivated by the need to retain strong visual representations while avoiding overfitting to perturbed or misleading multimodal inputs. ( Table 9 records the size for each dataset)
Detailed Experimental ResultsAppendix C): This section includes three key tables— Table 4,  Table 5 and  Table 6—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table 7 records the performance of different vanilla MLLMS under modality interference across modality-heavy datasets. We also include radar plots ( Figure 4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b [3] and InstructBlip-Vicuna-13b [2] and make further discussion on the selection of specific consistency loss.
Hyperparameter SettingsAppendix D): We present full training configurations used in our experiments, including optimization strategies, perturbation settings, and sampling ratios for different task types. This section enables reproducibility and highlights the computational efficiency of our proposed training scheme. We provide parameter analysis on iterations of PGD in  Figure 5.
Compute Resource DetailsAppendix E): We document hardware specifications, training durations, and resource costs for models of different scales. These details contextualize the feasibility of our approach in academic environments.
LimitationsAppendix F): We discuss the granularity of our current modality interference analysis, the selections of perturbations, and propose directions for more fine-grained future studies.
Broader ImpactsAppendix G): Finally, we reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
@@ Appendix C Detailed Experimental Results
Please see  Table 4,  Table 5,  Figure 4,  Table 7,  Table 6 for more details.
@@
Model Method ScienceQA-IMG MM-Bench-EN Seed-Bench-IMG VQA Overall
@@ C.1 Perturbation-based Evaluation Experiment Results
We conduct a controlled perturbation-based evaluation across various MLLMs, as shown in Table 7. Our results reveal that both vision and language tasks are vulnerable to cross-modal interference. In vision classification tasks, misleading textual descriptions (e.g., text contradicting image content) lead to severe performance drops. For example, InternVL2-2B and InstructBLIP-7B on Mini-ImageNet drop from 91.9% to 25.5% and from 92.0% to 13.6%, respectively. Conversely, for language tasks such as OpenBookQA and MMLU, irrelevant visual inputs—particularly semantically unrelated real images—also degrade performance. LLaVA-1.5-7B drops from 46.3% to 45.2% on MMLU, while InstructBLIP-13B sees over 5 points of degradation.
@@
To enable a more equitable comparison with existing multimodal models, we extend our method to two additional backbones: InstructBLIP-Vicuna-13B and QwenVL-2.5-7B. As shown in Table 6, our approach consistently improves performance across multiple VQA benchmarks, even under different instruction-tuning conditions, demonstrating its robustness and general applicability.
@@ Appendix D Hyper-Parameter Setting and Training details
Please see  Table 9 for more details.
@@ D.1 Parameter Analysis with PGD iterations
To investigate the effect of adversarial strength on model performance, we vary the number of PGD training iterations from 1 to 5 and evaluate the resulting VQA accuracy. As shown in  Figure 5, both LLaVA-1.5-7B and LLaVA-1.5-13B models benefit from PGD-based consistency training, with performance peaking at 2-step PGD (66.82% and 68.37%, respectively). Notably, excessive iterations (e.g., 4 or 5 steps) may lead to slight degradation, especially in larger models, likely due to over-perturbation and optimization difficulty.
@@ Appendix E Experiments compute resources
All experiments were conducted on 4×A6000 GPUs using DeepSpeed ZeRO-3 [40] with offloading. For 3B and 7B models, standard full fine-tuning takes approximately 8 and 10 hours, respectively. Adding KL regularization increases 7B training time to 15 hours, while introducing PGD-based perturbations further extends it to 18 hours (18 hours for 1 iteration, adding one iteration will add 2 hours). For larger-scale experiments with LLaVA-1.5-13B, we use 8×H100 GPUs with DeepSpeed ZeRO-2, where full fine-tuning completes in approximately 4 hours. We estimate the total compute cost to be within the typical academic scale.

### v2

@@ Appendix A Appendix Summary
Finetuning StrategiesAppendix˜B): We elaborate on our design choice to freeze the Q-Former in InstructBLIP-based models. This decision is motivated by the need to retain strong visual representations while avoiding overfitting to perturbed or misleading multimodal inputs. ( Table˜12 records the size for each dataset)
Detailed Experimental ResultsAppendix˜C): This section includes three key tables— Table˜4,  Table˜5 and  Table˜6—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table˜7 records the performance of different vanilla MLLMs under modality interference across modality-heavy datasets. We also include radar plots ( Figure˜4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b [3] and InstructBlip-Vicuna-13b [2] and make further discussion on the selection of specific consistency loss. In  Table˜8, we examine the generalization benefits of adversarial training by evaluating robustness under two types of out-of-distribution (OOD) perturbations: real-world OCR noise (from FUNSD [39]) and unrelated screenshots (from RICO [40]). In  Table˜9, we assess the impact of Chain-of-Thought prompting in mitigating modality interference, comparing its effectiveness against our method and standard baselines across both visual and textual modalities. In  Table˜10, we report results on the free-form generative VQA benchmark TextVQA [41], highlighting our method’s generalizability beyond multiple-choice formats.
Hyperparameter SettingsAppendix˜D): We present full training configurations used in our experiments, including optimization strategies, perturbation settings, and sampling ratios for different task types. This section enables reproducibility and highlights the computational efficiency of our proposed training scheme. We provide parameter analysis on iterations of PGD in  Figure˜5.
Compute Resource DetailsAppendix˜E): We document hardware specifications, training durations, and resource costs for models of different scales. These details contextualize the feasibility of our approach in academic environments.
LimitationsAppendix˜F): We discuss the granularity of our current modality interference analysis, the selections of perturbations, and propose directions for more fine-grained future studies.
Broader ImpactsAppendix˜G): Finally, we reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
@@ Appendix C Detailed Experimental Results
Please see  Table˜4,  Table˜5,  Figure˜4,  Table˜7,  Table˜6,  Table˜8,  Table˜9 and  Table˜10 for more details.
@@
Model Method ScienceQA-IMG $M$-Bench-EN Seed-Bench-IMG VQA Overall
@@ C.1 Perturbation-based Evaluation Experiment Results
We conduct a controlled perturbation-based evaluation across various MLLMs, as shown in Table˜7. Our results reveal that both vision and language tasks are vulnerable to cross-modal interference. In vision classification tasks, misleading textual descriptions (e.g., text contradicting image content) lead to severe performance drops. For example, InternVL2-2B and InstructBLIP-7B on Mini-ImageNet drop from 91.9% to 25.5% and from 92.0% to 13.6%, respectively. Conversely, for language tasks such as OpenBookQA and MMLU, irrelevant visual inputs—particularly semantically unrelated real images—also degrade performance. LLaVA-1.5-7B drops from 46.3% to 45.2% on MMLU, while InstructBLIP-13B sees over 5 points of degradation.
@@
To enable a more equitable comparison with existing multimodal models, we extend our method to two additional backbones: InstructBLIP-Vicuna-13B and QwenVL-2.5-7B. As shown in Table˜6, our approach consistently improves performance across multiple VQA benchmarks, even under different instruction-tuning conditions, demonstrating its robustness and general applicability.
@@ C.3 Discussion on KL&JS use for Consistency Regularization
C.4 Improving Generalizability through Adversarial Training
To assess the generalization benefits of adversarial training, we evaluate model robustness under two types of out-of-distribution (OOD) perturbations at test time:
  • Document OCR noise: Real-world noisy OCR snippets are sampled from the FUNSD dataset [39] and inserted as irrelevant textual distractors into visual classification tasks (Mini-ImageNet, Caltech-101).
  • Unrelated screenshots: Unrelated UI screenshots are drawn from the RICO dataset [40] and added as visual distractors to language-dominant VQA tasks (OpenBookQA, MMLU).
Each experiment is repeated multiple times, and we report average accuracy across runs (standard deviation << 0.1). Results are presented in Table 8. Across all model scales and task types, adversarial training consistently improves robustness to both types of perturbations. These findings indicate that the additional training overhead introduced by PGD is well-justified by the improved generalization to unseen distribution shifts—a desirable property for reliable deployment in real-world settings.
Table 8: Accuracy (%) on original and perturbed inputs. OCR snippets are inserted into image classification tasks, and RICO screenshots are added to VQA tasks.
Setting Model Mini-ImageNet Caltech-101 OpenBookQA MMLU
Origin OCR Origin OCR RandPixels Screenshot RandPixels Screenshot
vanilla instructblip-vicuna-7b 92.0 81.4 90.3 83.6 50.9 40.0 35.3 34.9
SFT instructblip-vicuna-7b 98.5 95.7 99.0 98.4 75.0 74.4 50.0 49.2
SFT + PGD instructblip-vicuna-7b 98.7 97.5 99.5 98.8 76.8 76.2 49.3 49.3
Ours instructblip-vicuna-7b 98.4 97.3 99.2 99.0 79.0 77.8 50.2 49.2
vanilla llava-1.5-7b 95.3 88.9 97.0 92.8 62.4 50.2 46.3 44.8
SFT llava-1.5-7b 98.2 98.0 98.5 98.0 78.6 78.0 51.1 50.6
SFT + PGD llava-1.5-7b 98.7 98.2 98.7 98.4 81.7 81.3 50.6 51.3
Ours llava-1.5-7b 98.6 98.2 99.3 99.0 81.8 81.2 51.5 51.3
vanilla qwen2.5-vl-7b 99.3 99.2 99.1 99.2 85.9 69.3 69.3 57.0
SFT qwen2.5-vl-7b 99.6 99.5 99.7 99.7 90.2 85.8 70.4 63.7
SFT + PGD qwen2.5-vl-7b 99.4 99.5 99.6 99.6 91.7 91.5 70.4 69.7
Ours qwen2.5-vl-7b 99.6 99.5 99.6 99.5 90.9 90.7 69.8 69.7
C.5 Evaluating Chain-of-Thought Prompting for Modality Interference Mitigation
To further investigate the potential of prompt-based methods in mitigating modality interference, we conduct additional experiments using Chain-of-Thought [44] style prompting. This approach aims to encourage structured reasoning by guiding the model through an explicit reasoning process before producing its final answer.
Specifically, we prepend the following CoT prompt to each input question:
Let’s think step by step:
   1. What information does the image provide?
   2. What is the question asking?
   3. Are there any misleading parts?
   4. Now give your final answer. Only write the final answer on a separate line like: ‘‘Answer: B’’
Model performance is evaluated under both clean inputs and perturbed conditions. For each task, we report accuracy on the original input as well as on the strongest-performing perturbation type (e.g., OCR insertion for visual classification tasks, or unrelated screenshots for VQA and language tasks). Results are presented in Table 9.
Table 9: Accuracy (%) under different interference settings across tasks and models. Each task includes original inputs and multiple types of perturbations.
Model Method Mini-ImageNet Caltech-101 OpenBookQA MMLU
Orig UF MD Orig UF MD RP RI FB FW RP RI FB FW
vanilla llava-1.5-7b 95.3 93.4 43.5 97.0 95.9 57.4 62.4 56.4 62.5 63.4 46.3 45.2 45.9 45.8
CoT llava-1.5-7b 81.7 70.9 28.9 80.9 71.8 36.0 38.8 38.9 41.0 41.0 39.6 38.7 40.2 40.4
Ours llava-1.5-7b 98.6 98.4 98.7 99.3 98.9 99.3 81.8 81.0 81.7 81.7 51.5 50.9 51.5 51.4
vanilla llava-1.5-13b 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
CoT llava-1.5-13b 92.9 85.9 62.8 96.6 92.1 67.8 55.6 53.0 56.5 57.3 47.3 45.5 47.3 47.8
Ours llava-1.5-13b 98.7 97.9 98.0 98.7 98.7 98.8 83.2 81.2 83.8 83.0 52.2 51.6 52.3 53.4
vanilla qwen2.5-vl-3b 98.9 98.5 94.9 98.8 99.0 94.4 79.9 74.6 80.0 79.7 63.5 61.1 64.0 63.6
CoT qwen2.5-vl-3b 92.3 96.8 88.2 94.7 96.1 86.0 61.0 52.8 61.3 61.1 51.5 48.7 51.6 51.4
Ours qwen2.5-vl-3b 99.3 99.2 99.2 99.7 99.7 99.5 86.7 86.4 86.6 86.6 64.8 64.5 65.0 65.1
vanilla qwen2.5-vl-7b 99.3 99.3 96.3 99.1 98.9 97.2 85.9 77.5 85.8 86.0 69.3 63.7 68.9 68.9
CoT qwen2.5-vl-7b 99.1 98.9 91.4 98.5 98.5 92.5 77.8 68.8 77.3 77.6 61.5 57.4 60.8 60.9
Ours qwen2.5-vl-7b 99.6 99.5 99.5 99.6 99.6 99.7 90.9 89.7 90.8 91.0 69.8 68.2 69.8 70.0
C.6 Evaluation on Free-Form VQA
To further assess the generalizability of our method beyond multiple-choice VQA tasks, we evaluate on TextVQA [41], a free-form generative visual question answering dataset that requires reasoning over both textual and visual content in natural images. We follow the benchmark adopted by LLaVA [1], which evaluates a model’s ability to both recognize textual characters within images and effectively handle noisy outputs generated by OCR systems.
Following standard evaluation protocols and existing MLLM baselines (e.g., LLaVA, Qwen2.5-VL, InstructBLIP), we report model performance averaged over multiple runs (standard deviation << 0.4). Results are presented in Table 10.
Our method achieves consistent improvements across most model families, indicating its effectiveness not only in MCQA scenarios but also in open-ended multimodal reasoning settings.
Table 10: Accuracy (%) on the TextVQA dataset across different model families.
Setting Model TextVQA
Vanilla InstructBLIP-Vicuna-7B 19.7
Ours InstructBLIP-Vicuna-7B 32.4
Vanilla LLaVA-1.5-7B 49.8
Ours LLaVA-1.5-7B 51.2
Vanilla Qwen2.5-VL-3B 77.4
Ours Qwen2.5-VL-3B 75.3
Vanilla Qwen2.5-VL-7B 81.4
Ours Qwen2.5-VL-7B 84.8
@@ Appendix D Hyper-Parameter Setting and Training details
Please see  Table˜12 for more details.
@@ D.1 Parameter Analysis with PGD iterations
To investigate the effect of adversarial strength on model performance, we vary the number of PGD training iterations from 1 to 5 and evaluate the resulting VQA accuracy. As shown in  Figure˜5, both LLaVA-1.5-7B and LLaVA-1.5-13B models benefit from PGD-based consistency training, with performance peaking at 2-step PGD (66.82% and 68.37%, respectively). Notably, excessive iterations (e.g., 4 or 5 steps) may lead to slight degradation, especially in larger models, likely due to over-perturbation and optimization difficulty.
@@ Appendix E Experiments compute resources
[…]
To quantify the computational overhead introduced by our PGD-based training, we provide both theoretical FLOPs analysis and empirical wall-clock training time across model scales.
In the standard supervised fine-tuning (SFT) setting, the FLOPs per batch can be approximated as:
$\text{FLOPs}_{\text{SFT}}\approx B_{s}\cdot(f_{\text{LLM}}+b_{\text{LLM}}),$ (11)
where $B_{s}$ is the batch size, $f_{\text{LLM}}$ denotes the FLOPs of a forward pass through the LLM, and $b_{\text{LLM}}$ the backward pass. In our PGD-enhanced training, each sample undergoes $N$ PGD steps, each requiring an additional forward pass through the frozen LLM. Since gradients are computed only with respect to input embeddings via torch.autograd.grad, the overhead is minimal and thus ignored. After perturbation, clean and adversarial inputs are concatenated, resulting in a forward cost of $2B_{s}\cdot f_{\text{LLM}}$, followed by one backward pass. The total cost becomes:
$\text{FLOPs}_{\text{PGD+SFT}}=B_{s}\cdot(Nf_{\text{LLM}}+2f_{\text{LLM}}+b_{\text{LLM}}).$ (12)
The relative overhead compared to vanilla SFT is:
$\frac{Nf+2f+b}{f+b}.$ (13)
Assuming $b_{\text{LLM}}\approx 2f_{\text{LLM}}$, this simplifies to:
$\frac{N+4}{3}.$ (14)
For our default $N=1$, the theoretical FLOPs increase to approximately $1.66\times$ that of SFT.
We further report the actual training time across model scales, as shown in Table 13.
Table 13: Training time (in hours) across model scales. $\Delta$ denotes additional overhead per PGD iteration step.
Model SFT SFT + KL SFT + PGD Ours (PGD + KL)
Qwen2.5-VL-3B 1.5h 1.5–1.75h 3–3.5h ($\Delta$ = 0.5h) 3.5–4h ($\Delta$ = 0.5h)
LLaVA-1.5-7B 4h 4–4.5h 6–6.5h ($\Delta$ = 0.5h) 6–6.5h ($\Delta$ = 0.5h)
LLaVA-1.5-13B 10h 10–11h 13–14h ($\Delta$ = 1h) 13–14h ($\Delta$ = 1h)
Although the theoretical FLOPs suggest a $\sim$66% increase in cost when $N=1$, the actual wall-clock time increase is much smaller. This is because our PGD implementation leverages forward-only passes over frozen LLMs, avoiding costly backward and optimizer updates. As a result, the added runtime remains modest even on large models (e.g., only +2.5h for LLaVA-13B). Moreover, KL consistency training introduces negligible overhead compared to SFT.